OpenStudy (anonymous):

How do I graph 3 > or equal to x - 3y? Please help me out!!!

OpenStudy (anonymous):

Well you have to get y on one side by itself so \[3 \ge x-3y\] Subtract x on both sides \[-x +3 \ge -3y\] Then divide by -3 on both sides. When you divide by a negative you flip the sign so in the case it goes from greater than or equal to to less then or equal to. \[\frac{ 1 }{ 3 }x -1 \le y\] It could also be written like this. \[y \ge \frac{ 1 }{ 3 }x -1\]

OpenStudy (anonymous):

You can use this site to graph. http://www.garrettbartley.com/graphpaper.html

OpenStudy (anonymous):

Thanx but can u show me how to do it?

OpenStudy (anonymous):

To determine which half to shade you just plug in a point. I'm going to use (0,0) \[0 \ge \frac{ 1 }{ 3 }(0)-1\] \[0 \ge 0 -1\] \[0 \ge -1\] Since this is true shade the half that has the point (0,0)

OpenStudy (anonymous):

What are the x and y I mean wat are the points?

OpenStudy (anonymous):

Like the x and y intercepts?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The the y-int is given to you in the inequality you have y "=" mx+b where b is the y-int So you y-int is -1 So where does the line hit the x-axis? (The x-axis is the horizontal black line)

OpenStudy (anonymous):

No clue

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Just look at my attachment and look where the orange line hits the black horizontal line?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yepp.

OpenStudy (anonymous):

Yaaay!

OpenStudy (anonymous):

I think I got how to do it byt y did u do it as slope intercept form? Y didnt u do it normally

OpenStudy (anonymous):

That is normally. That's how you graph.

OpenStudy (anonymous):

K thanx